5 - X/3 = 3(X+1)/7
Please and Thanks!

Answer by onesavedall(1)

Np,
5-(X/3)=3(x+1)/7
[5-(X/3)=3(x+1)/7]21
105-7x=9(x+1)
105-7x=9x+9
subtract 9 from both sides
96-7x= 9x
add 7x to both sides
(96=12x)/12
x=96/12
x=8